# Context-Sensitive Abstractions for Reinforcement Learning with Parameterized Actions

### v1

@@ 1 Introduction
Reinforcement learning (RL) has delivered strong results across a diverse range of decision-making tasks, from discrete action settings like Atari games (mnih2015human) to continuous control scenarios such as robotic manipulation (schulman2017proximal). Yet most leading RL approaches (schulman2017proximal; haarnoja2018soft; schrittwieser2020mastering; hansen2023td) are designed for either discrete or continuous action spaces—not both. Many real-world problems violate this dichotomy. In autonomous driving, for example, the agent must choose among qualitatively distinct actions (accelerate, brake, turn), each endowed with discrete or continuous parameters such as braking force or steering angle. Such actions—known as parameterized actions—require choosing not only the action but also determine its (real-valued) parameters before execution.
@@
While recent methods have made progress in addressing parameterized actions  (xiong2018parametrized; bester2019multi; li2021hyar), they largely ignore utilizing the underlying structure inherent in parameterized-action spaces. In navigation tasks, for instance, an agent should adjust movement parameters with high precision near obstacles but can act with much coarser control in open areas. Existing approaches also often rely on carefully engineered dense rewards and environment-specific initializations to facilitate learning or benefit from relatively short “effective horizons” to remain tractable (laidlaw2023bridging). A detailed discussion of related work is in Sec. 5.
@@
Given an input problem in the RL setting where a state is expressed using discrete and continuous state variables and an action is expressed using continuous or discrete parameters, PEARL learns context-sensitive abstractions while performing $TD(\lambda)$. It uses a combination of dispersion in TD-error and value-function signals to learn which abstract states and action parameters require finer resolution during learning. Our approach builds upon our recent work on conditional state abstractions (dadvar2023conditional), and introduces new algorithms for learning more general forms of state abstractions alongside abstractions of action parameters.
@@ 2 Preliminaries
We use the framework of episodic factored goal-oriented Markov decision process (MDP) with parameterized actions (bertsekas2011dynamic; hausknecht2015deep; deng2022polynomial). An MDP $\mathcal{M}$} is defined as $\langle{\mathcal{V}},{\mathcal{S}},{\mathcal{A}},T,R,\gamma,h,s_{0},G\rangle$, where $\mathcal{V}$} is a set of state variables and the domain of each variable $v\in{\mathcal{V}}$ is a bounded interval ${\mathcal{D}}_{v_{i}}=\left[{\mathcal{D}}^{\emph{min}}_{v_{i}},{\mathcal{D}}^{\emph{max}}_{v_{i}}\right]\subseteq\mathbb{R}$; $\mathcal{S}$} denotes the set of factored states defined by $\mathcal{V}$}, where a state $s\in{\mathcal{S}}$ is an assignment of values to all variables in $\mathcal{V}$}: $s=\{v_{i}=x_{k}|v_{i}\in{\mathcal{V}}\land x_{k}\in{\mathcal{D}}_{v_{i}}\}$. We use $s(v_{i})$ to denote the value of variable $v_{i}$ in state $s$.
@@
The objective is to learn a policy $\pi_{\mathcal{M}}:{\mathcal{S}}\rightarrow\tilde{{\mathcal{A}}}$ that when executed from the initial state $s_{0}$, reaches a goal state in $s_{g}\in G$ while maximizing the expected cumulative discounted reward $\mathbb{E}_{\pi}[\sum_{t=0}^{t=h}\gamma^{t}r_{t}]$. We use the RL setting, where both $T$ and $R$ are unknown (sutton1998introduction).
@@
We use TD($\lambda$(sutton1988learning) for learning the policy $\pi$. It combines one-step TD and Monte Carlo methods by weighting updates across multiple future time steps, controlled by the trace-decay parameter $\lambda\in[0,1]$.
@@
Abstraction has been recognized as a key mechanism for achieving scalability in long horizon, sparse reward settings  (li2006towards; shah2024hierarchical; wang2024building). A state abstraction is a mapping $\alpha:\mathcal{S}\rightarrow\overline{\mathcal{S}}$ that assigns each concrete state $s\in\mathcal{S}$ to an abstract state $\overline{s}\in\overline{\mathcal{S}}$, where $\overline{\mathcal{S}}$ is a partitioning of the original state space $\mathcal{S}$. In this work, we define an analogous notion of abstraction for an action, defined as a partitioning of the action-parameter space (formalized in Sec. 3.1).
@@ 3 Our Approach
Consider an AI agent in an Office environment (Fig. 1) that must collect and deliver a coffee and a mail between rooms and offices. The state variables include the agent’s $(x,y)$ position with $x,y\in[0.0,5.0)$, and two binary variables: $c\in\{0,1\}$ and $m\in\{0,1\}$ indicating whether it is carrying coffee or mail. The agent has four actions to move in the cardinal directions, i.e., ${\mathcal{A}}=\{\emph{up}(d),\emph{down}(d),\emph{left}(d),\emph{right}(d)\}$, each with one continuous parameter $d\in[0,0.5)$ that determines the movement distance. Actions may result in stochastic displacements along orthogonal directions, and the agent picks or drops items automatically at designated locations. This setting extends the OfficeWorld environment  (icarte2022reward) by incorporating parameterized actions.
@@
This work builds upon our prior work (dadvar2023conditional) which learned state abstractions with strictly uniform refinements. It did not address the problem of parameterized actions and offered no mechanism for more flexible refinements of abstractions. The abstraction learned in this work has the following desirable properties: (i) The abstractions are flexible—the abstract state boundaries are not constrained to be orthogonal or axis-aligned. This flexibility allows the learned abstractions to better adapt to the geometry and dynamics of the environment—for example, by placing boundaries where agent behavior changes, following the contours of obstacles. Fig. 1 illustrates an example of such a state abstraction for Office World, where each colored region represents a distinct abstract state. (ii) Moreover, each abstract state has a conjoined action parameter tree for each action (shown in Fig. 2), allowing varying levels of precision in different abstract states. E.g., in open areas—such as the centers of rooms (e.g., abstract state $\overline{s}_{7}$)—the agent can move freely without requiring high precision in selecting movement distances (e.g., abstract action left([0.25,0.5))). In contrast, in more constrained areas—such as corridors, near obstacles, or narrow passages (e.g., abstract state $\overline{s}_{8}$)—precise control over movement is crucial (e.g., abstract action left([0.0,0.1))). Fig. 2 shows the unified state-action abstraction tree, where the leaves represent these abstract states and abstract actions. The shown abstractions capture the required higher precision for selecting action parameters. We hypothesize that automatically identifying such meaningful abstract states and corresponding action parameter trees can significantly improve sample efficiency in policy learning.
@@ 3.2 Learning Abstraction Trees
It is well-known that value-function is inaccurate at start of Q-learning, thus unsuitable for early refinement, whereas, TD-error better represents similar futures during early learning (e.g., (kearns1998finite)). Thus, in early stages of learning, we rely on the value of the temporal difference (TD) error to provide a stronger signal of behavioral inconsistency: if $\delta(\bar{s},\bar{a})$ values show a high standard deviation ($SD$) for an abstract state-action pair in $\bar{{\mathcal{D}}}$, then this abstract pair may be representing heterogeneous regions where the $Q$ function is changing at significantly different rates. On the other hand, as learning progresses and the policy stabilizes, value function estimates become more reliable. At this point the variability of value-function estimates across concrete states in an abstract state are a better indicator of heterogeneity in the abstraction. Therefore, we blend TD error and value-function dispersion metrics. Since it is infeasible to maintain a tabular representation (e.g., a Q-table) over all continuous concrete states and actions, we compute an estimate of the concrete-state value function as follows:
$\hat{V}(s_{i})=r_{i}+\gamma\max_{\overline{a}}\overline{Q}(\overline{s}_{i+1},\overline{a})-Q(s_{i},\overline{a})$ (2)
@@
Here, the standard deviation is computed over all occurrences of the pair $\bar{s}_{i},\bar{a}_{i}$ in $\bar{{\mathcal{D}}}$. A scheduling mechanism is used to gradually shift emphasis from TD error dispersions to value function dispersions. This is achieved with a weighting parameter $\beta$, initialized at 1.0 and annealed across episodes by a decay schedule. States with high heterogeneity, under the current $\beta$, are selected for refinement into finer abstractions.
@@
Multiple paradigms can be used for refining the abstract state-action regions that feature a high heterogeneity under this formulation. We consider two paradigms: uniform refinement as proposed in prior work (dadvar2023conditional), and a novel flexible refinement that uses statistical learning. Both state abstractions (nodes for SPA-CATs) and action abstractions (nodes for Action Trees) can be refined using these methods. However, for brevity, we describe them in the context of refining state abstractions.
@@
Given an abstract state $\overline{s}$ selected for refinement, uniform partitioning bisects the interval corresponding to each (variable) independently, resulting in an orthogonal binary tree decomposition of the state space (see Fig. 3 (right)). While straightforward, such abstractions are best suited for domains where the Q-function can be factorized into functions over individual state variables and they require extensive refinements to express regions that feature homogeneous value function estimates, but do not constitute hypercubes.
@@
We introduce a novel learning-based approach for constructing flexible refinements. Given an abstract state $\overline{s}$ selected for refinement and the associated set of execution traces, we partition $\overline{s}$ into at most $\mathcal{K}$} finer abstract states by clustering the concrete states contained within $\overline{s}$. Specifically, we apply Agglomerative Clustering (murtagh2012algorithms) from scikit-learn (pedregosa2011scikit) with an adaptive distance threshold: starting from 0.1, we incrementally increase the threshold by 0.001 until the number of clusters is below a specified maximum. This prevents over-fragmentation while ensuring meaningful behavioral distinctions are captured. We use the following similarity criterion to form coherent partitions that reflect underlying behavioral distinctions:
@@ 3.3 PEARL Algorithm
In this phase, the agent learns an abstract policy $\pi:\overline{\mathcal{S}}\rightarrow\overline{\mathcal{A}}$ over the current SPA-CAT structure using tabular TD-$\lambda$ (sutton1988learning) for $n_{\emph{refine}}$ episodes (lines 4–11). During each episode, the agent follows the abstract policy by executing the corresponding abstract action in the current abstract state, continuing until it reaches a new abstract state or the episode terminates (lines 6–7).
@@ 4 Empirical Results
We evaluate on domains well-established as challenging (illustrated in Fig.  4): (i) OfficeWorld (icarte2022reward; corazza2024expediting) (ii) Pinball (roice2024new; rodriguez2024learning), (iii) Multi-city transport (ma2021hierarchical; oswald2024large), and (iv) Robot Soccer Goal (bester2019multi). Among these, former three are especially challenging due to longer effective planning horizons.
@@
Standard RL approaches—tabular RL (sutton1988learning; watkins1989learning), deep RL (mnih2015human; lillicrap2015continuous; schulman2017proximal; haarnoja2018soft), hierarchical RL (nachum2018data; levy2017learning)—are not designed to handle parameterized actions, making them unsuitable as baselines. We therefore compare PEARL against two baselines that support parameterized actions: $(i)$ MP-DQN (bester2019multi), which extends P-DQN (xiong2018parametrized) by combining DQN and DDPG while addressing P-DQN’s over-parameterization problem through multi-pass processing, and $(ii)$ HyAR (li2021hyar), which learns latent space of hybrid action space and models dependencies between discrete action and continuous parameter using an embedding table and a conditional Variational Auto-Encoder (VAE). To evaluate and compare learning performance fairly without manually biasing learning with head-starts towards favorable solutions, we used the original source code for these baselines while removing their hand-crafted, environment-specific weight initializations. We replaced them with zero or randomized initializations, whichever yielded better performance. This provides a consistent, unbiased evaluation of each method’s true learning capability.
@@ 5 Related Work
Most standard RL methods  (mnih2015human; lillicrap2015continuous; schulman2017proximal; haarnoja2018soft) are designed for homogeneous action spaces, handling either purely discrete or purely continuous action spaces. Moreover, their success has mostly been limited to settings with short effective horizons, where multi-step lookahead is unnecessary (laidlaw2023bridging). Parameterized actions, which combine discrete actions with associated continuous parameters, present additional challenges they do not address. Some early approaches, such as Q-PAMDP (masson2016reinforcement) alternate between optimizing discrete actions and their continuous parameters. PADDPG (hausknecht2015deep) collapses all action parameters into a single continuous vector. These methods do not exploit the inherent structure of the parameterized actions (the dependency between discrete actions and their associated parameters) essential for learning effective policies.
P-DQN (xiong2018parametrized) directly handles hybrid action spaces without relaxation or approximation by integrating a DQN (to deal with discrete actions) and a DDPG (to deal with continuous actions). However, this approach treats all action-parameters as a single joint input to the Q-network, which results in dependence of each discrete action’s value on all action-parameters, not only those associated with that action. To overcome the over-parameterization problem of P-DQN, MP-DQN (bester2019multi) extend P-DQN with a multiple-pass mechanism, splitting the action-parameter inputs to the Q-network using several passes. H-PPO (fan2019hybrid) decomposes the action space using parallel sub-actor networks—one for discrete action selection and others for parameter learning—guided by a shared critic. HyAR (li2021hyar) learns a latent representation for hybrid actions via a variational autoencoder, enabling standard DRL algorithms. These methods incur added computational cost due to architectural complexity and hyperparameter sensitivity.
@@
Coarse-to-fine RL (CRL) (seo2024continuous) discretize continuous action spaces by learning a single action discretization that spans the entire state space. This is achieved by independently learning a Q-network for each action dimension. In contrast, our method learns distinct abstractions of parameterized actions conditioned on abstract states. Unlike prior top-down abstraction methods limited to discrete actions (dadvar2023conditional; nayyar2025autonomous), PEARL handles parameterized actions with continuous parameters via action abstraction and supports flexible refinement to compactly capture structure in the problem.

### v2

@@ 1 Introduction
Reinforcement learning (RL) has delivered strong results across a diverse range of decision-making tasks, from discrete action settings like Atari games (Mnih et al. 2015) to continuous control scenarios such as robotic manipulation (Schulman et al. 2017). Yet most leading RL approaches (Schulman et al. 2017; Haarnoja et al. 2018; Schrittwieser et al. 2020; Hansen et al. 2024) are designed for either discrete or continuous action spaces—not both. Many real-world problems violate this dichotomy. In autonomous driving, for example, the agent must choose among qualitatively distinct actions (accelerate, brake, turn), each endowed with discrete or continuous parameters such as braking force or steering angle. Such actions—known as parameterized actions—require choosing not only the action but also determine its (real-valued) parameters before execution.
@@
While recent methods have made progress in addressing parameterized actions  (Xiong et al. 2018; Bester et al. 2019; Li et al. 2022), they largely ignore utilizing the underlying structure inherent in parameterized-action spaces. In navigation tasks, for instance, an agent should adjust movement parameters with high precision near obstacles but can act with much coarser control in open areas. Existing approaches also often rely on carefully engineered dense rewards and environment-specific initializations to facilitate learning or benefit from relatively short “effective horizons” to remain tractable (Laidlaw et al. 2023). A detailed discussion of related work is in Sec. 5.
@@
Given an input problem in the RL setting where a state is expressed using discrete and continuous state variables and an action is expressed using continuous or discrete parameters, PEARL learns context-sensitive abstractions while performing $TD(\lambda)$. It uses a combination of dispersion in TD-error and value-function signals to learn which abstract states and action parameters require finer resolution during learning. Our approach builds upon our recent work on conditional state abstractions (Dadvar et al. 2023), and introduces new algorithms for learning more general forms of state abstractions alongside abstractions of action parameters.
@@ 2 Preliminaries
We use the framework of episodic factored goal-oriented Markov decision process (MDP) with parameterized actions (Bertsekas and others 2011; Hausknecht and Stone 2016; Deng et al. 2022). An MDP $\mathcal{M}$} is defined as $\langle{\mathcal{V}},{\mathcal{S}},{\mathcal{A}},T,R,\gamma,h,s_{0},G\rangle$, where $\mathcal{V}$} is a set of state variables and the domain of each variable $v\in{\mathcal{V}}$ is a bounded interval ${\mathcal{D}}_{v_{i}}=\left[{\mathcal{D}}^{\emph{min}}_{v_{i}},{\mathcal{D}}^{\emph{max}}_{v_{i}}\right]\subseteq\mathbb{R}$; $\mathcal{S}$} denotes the set of factored states defined by $\mathcal{V}$}, where a state $s\in{\mathcal{S}}$ is an assignment of values to all variables in $\mathcal{V}$}: $s=\{v_{i}=x_{k}|v_{i}\in{\mathcal{V}}\land x_{k}\in{\mathcal{D}}_{v_{i}}\}$. We use $s(v_{i})$ to denote the value of variable $v_{i}$ in state $s$.
@@
The objective is to learn a policy $\pi_{\mathcal{M}}:{\mathcal{S}}\rightarrow\tilde{{\mathcal{A}}}$ that when executed from the initial state $s_{0}$, reaches a goal state in $s_{g}\in G$ while maximizing the expected cumulative discounted reward $\mathbb{E}_{\pi}[\sum_{t=0}^{t=h}\gamma^{t}r_{t}]$. We use the RL setting, where both $T$ and $R$ are unknown (Sutton and Barto 1998).
@@
We use TD($\lambda$(Sutton 1988) for learning the policy $\pi$. It combines one-step TD and Monte Carlo methods by weighting updates across multiple future time steps, controlled by the trace-decay parameter $\lambda\in[0,1]$.
@@
Abstraction has been recognized as a key mechanism for achieving scalability in long horizon, sparse reward settings  (Li et al. 2006; Shah and Srivastava 2024; Wang et al. 2024). A state abstraction is a mapping $\alpha:\mathcal{S}\rightarrow\overline{\mathcal{S}}$ that assigns each concrete state $s\in\mathcal{S}$ to an abstract state $\overline{s}\in\overline{\mathcal{S}}$, where $\overline{\mathcal{S}}$ is a partitioning of the original state space $\mathcal{S}$. In this work, we define an analogous notion of abstraction for an action, defined as a partitioning of the action-parameter space (formalized in Sec. 3.1).
@@ 3 Our Approach
Consider an AI agent in an Office environment (Fig. 1) that must collect and deliver a coffee and a mail between rooms and offices. The state variables include the agent’s $(x,y)$ position with $x,y\in[0.0,5.0)$, and two binary variables: $c\in\{0,1\}$ and $m\in\{0,1\}$ indicating whether it is carrying coffee or mail. The agent has four actions to move in the cardinal directions, i.e., ${\mathcal{A}}=\{\emph{up}(d),\emph{down}(d),\emph{left}(d),\emph{right}(d)\}$, each with one continuous parameter $d\in[0,0.5)$ that determines the movement distance. Actions may result in stochastic displacements along orthogonal directions, and the agent picks or drops items automatically at designated locations. This setting extends the OfficeWorld environment  (Icarte et al. 2022) by incorporating parameterized actions.
@@
This work builds upon our prior work (Dadvar et al. 2023) which learned state abstractions with strictly uniform refinements. It did not address the problem of parameterized actions and offered no mechanism for more flexible refinements of abstractions. The abstraction learned in this work has the following desirable properties: (i) The abstractions are flexible—the abstract state boundaries are not constrained to be orthogonal or axis-aligned. This flexibility allows the learned abstractions to better adapt to the geometry and dynamics of the environment—for example, by placing boundaries where agent behavior changes, following the contours of obstacles. Fig. 1 illustrates an example of such a state abstraction for Office World, where each colored region represents a distinct abstract state. (ii) Moreover, each abstract state has a conjoined action parameter tree for each action (shown in Fig. 2), allowing varying levels of precision in different abstract states. E.g., in open areas—such as the centers of rooms (e.g., abstract state $\overline{s}_{7}$)—the agent can move freely without requiring high precision in selecting movement distances (e.g., abstract action left([0.25,0.5))). In contrast, in more constrained areas—such as corridors, near obstacles, or narrow passages (e.g., abstract state $\overline{s}_{8}$)—precise control over movement is crucial (e.g., abstract action left([0.0,0.1))). Fig. 2 shows the unified state-action abstraction tree, where the leaves represent these abstract states and abstract actions. The shown abstractions capture the required higher precision for selecting action parameters. We hypothesize that automatically identifying such meaningful abstract states and corresponding action parameter trees can significantly improve sample efficiency in policy learning.
@@ 3.2 Learning Abstraction Trees
It is well-known that value-function is inaccurate at start of Q-learning, thus unsuitable for early refinement, whereas, TD-error better represents similar futures during early learning (e.g., (Kearns and Singh 1998)). Thus, in early stages of learning, we rely on the value of the temporal difference (TD) error to provide a stronger signal of behavioral inconsistency: if $\delta(\bar{s},\bar{a})$ values show a high standard deviation ($SD$) for an abstract state-action pair in $\bar{{\mathcal{D}}}$, then this abstract pair may be representing heterogeneous regions where the $Q$ function is changing at significantly different rates. On the other hand, as learning progresses and the policy stabilizes, value function estimates become more reliable. At this point the variability of value-function estimates across concrete states in an abstract state are a better indicator of heterogeneity in the abstraction. Therefore, we blend TD error and value-function dispersion metrics. Since it is infeasible to maintain a tabular representation (e.g., a Q-table) over all continuous concrete states and actions, we compute an estimate of the concrete-state value function as follows:
$\hat{\delta}(s_{i},\overline{a}_{i})$ $=r_{i}+\gamma\max_{\overline{a}}\overline{Q}(\overline{s}_{i+1},\overline{a})-Q(s_{i},\overline{a}_{i})$ (2)
$\hat{V}(s_{i})$ $=\max_{\overline{a}}[Q(s_{i},\overline{a})+\alpha*\hat{\delta}(s_{i},\overline{a})]$
@@
Here, the standard deviation is computed over all occurrences of the pair $\bar{s}_{i},\bar{a}_{i}$ in $\bar{{\mathcal{D}}}$. A scheduling mechanism is used to gradually shift emphasis from TD error to value function dispersions. This is achieved with a weighting parameter $\beta$, initialized at 1.0 and annealed across episodes by a decay schedule. States with high heterogeneity, under the current $\beta$, are selected for refinement into finer abstractions.
@@
Multiple paradigms can be used for refining the abstract state-action regions that feature a high heterogeneity under this formulation. We consider two paradigms: uniform refinement as proposed in prior work (Dadvar et al. 2023), and a novel flexible refinement that uses statistical learning. Both state abstractions (nodes for SPA-CATs) and action abstractions (nodes for Action Trees) can be refined using these methods. However, for brevity, we describe them in the context of refining state abstractions.
@@
Uniform partitioning refines an abstract state $\overline{s}$ by independently bisecting each variable’s interval to create an orthogonal binary tree decomposition of the state space (see Fig. 3 (right)). While straightforward, such abstractions are best suited for domains where the Q-function can be factorized into functions over individual state variables and they require extensive refinements to express regions that feature homogeneous value function estimates, but do not constitute hypercubes.
@@
We introduce a novel learning-based approach for constructing flexible refinements. Given an abstract state $\overline{s}$ selected for refinement and the associated set of execution traces, we partition $\overline{s}$ into at most $\mathcal{K}$} finer abstract states by clustering the concrete states contained within $\overline{s}$. Specifically, we apply Agglomerative Clustering (Murtagh and Contreras 2012) from scikit-learn (Pedregosa et al. 2011) with an adaptive distance threshold: starting from 0.1, we incrementally increase the threshold by 0.001 until the number of clusters is below a specified maximum. This prevents over-fragmentation while ensuring meaningful behavioral distinctions are captured. We use the following similarity criterion to form coherent partitions that reflect underlying behavioral distinctions:
@@ 3.3 PEARL Algorithm
In this phase, the agent learns an abstract policy $\pi:\overline{\mathcal{S}}\rightarrow\overline{\mathcal{A}}$ over the current SPA-CAT structure using tabular TD-$\lambda$ (Sutton 1988) for $n_{\emph{refine}}$ episodes (lines 4–11). During each episode, the agent follows the abstract policy by executing the corresponding abstract action in the current abstract state, continuing until it reaches a new abstract state or the episode terminates (lines 6–7).
@@ 4 Empirical Results
We evaluate on domains well-established as challenging (illustrated in Fig.  4): (i) OfficeWorld (Icarte et al. 2022; Corazza et al. 2024) (ii) Pinball (Roice et al. 2024; Rodriguez-Sanchez and Konidaris 2024), (iii) Multi-city transport (Ma et al. 2021; Oswald et al. 2024), and (iv) Robot Soccer Goal (Bester et al. 2019). Among these, former three are especially challenging due to longer effective planning horizons.
@@
Standard RL approaches—tabular RL (Sutton 1988; Watkins and others 1989), deep RL (Mnih et al. 2015; Lillicrap et al. 2015; Schulman et al. 2017; Haarnoja et al. 2018), hierarchical RL (Nachum et al. 2018; Levy et al. 2019)—are not designed to handle parameterized actions, making them unsuitable as baselines. We therefore compare PEARL against two baselines that support parameterized actions: $(i)$ MP-DQN (Bester et al. 2019), which extends P-DQN (Xiong et al. 2018) by combining DQN and DDPG while addressing P-DQN’s over-parameterization problem through multi-pass processing, and $(ii)$ HyAR (Li et al. 2022), which learns latent space of hybrid action space and models dependencies between discrete action and continuous parameter using an embedding table and a conditional Variational Auto-Encoder (VAE). To evaluate and compare learning performance fairly without manually biasing learning with head-starts towards favorable solutions, we used the original source code for these baselines while removing their hand-crafted, environment-specific weight initializations. We replaced them with zero or randomized initializations, whichever yielded better performance. This provides a consistent, unbiased evaluation of each method’s true learning capability.
@@ 5 Related Work
Most standard RL methods  (Mnih et al. 2015; Lillicrap et al. 2015; Schulman et al. 2017; Haarnoja et al. 2018) are designed for homogeneous action spaces, handling either purely discrete or purely continuous action spaces. Moreover, their success has mostly been limited to settings with short effective horizons, where multi-step lookahead is unnecessary (Laidlaw et al. 2023). Parameterized actions, which combine discrete actions with associated continuous parameters, present additional challenges they do not address. Some early approaches, such as Q-PAMDP (Masson et al. 2016) alternate between optimizing discrete actions and their continuous parameters. PADDPG (Hausknecht and Stone 2016) collapses all action parameters into a single continuous vector. These methods do not exploit the inherent structure of the parameterized actions (the dependency between discrete actions and their associated parameters) essential for learning effective policies.
P-DQN (Xiong et al. 2018) directly handles hybrid action spaces without relaxation or approximation by integrating a DQN (to deal with discrete actions) and a DDPG (to deal with continuous actions). However, this approach treats all action-parameters as a single joint input to the Q-network, which results in dependence of each discrete action’s value on all action-parameters, not only those associated with that action. To overcome the over-parameterization problem of P-DQN, MP-DQN (Bester et al. 2019) extend P-DQN with a multiple-pass mechanism, splitting the action-parameter inputs to the Q-network using several passes. H-PPO (Fan et al. 2019) decomposes the action space using parallel sub-actor networks—one for discrete action selection and others for parameter learning—guided by a shared critic. HyAR (Li et al. 2022) learns a latent representation for hybrid actions via a variational autoencoder, enabling standard DRL algorithms. These methods incur added computational cost due to architectural complexity and hyperparameter sensitivity.
@@
Coarse-to-fine RL (CRL) (Seo et al. 2025) discretize continuous action spaces by learning a single action discretization that spans the entire state space. This is achieved by independently learning a Q-network for each action dimension. In contrast, our method learns distinct abstractions of parameterized actions conditioned on abstract states. Unlike prior top-down abstraction methods limited to discrete actions (Dadvar et al. 2023; Nayyar and Srivastava 2025), PEARL handles parameterized actions with continuous parameters via action abstraction and supports flexible refinement to compactly capture structure in the problem.